OpenStudy (amenah8):

Find the indeterminate form and locate the limit using l'hopital's rule: Find the limit as x approaches 1 at (x^2 - 2x + 1)^x-1

OpenStudy (amenah8):

i'm going to bed now :) but feel free to help me through the problem! i will be back before my class at 11:00am :). I will medal!!!!

OpenStudy (freckles):

so the problem is \[\lim_{x \rightarrow 1}[(x^2-2x+1)^{x}-1 ]\]?

OpenStudy (freckles):

for some reason I think the question is to evaluate: \[\lim_{x \rightarrow 1^+} (x^2-2x+1)^{x-1}\] I would write x^2-2x+1 as (x-1)^2 and then I would use e^(ln(y))=y somewhere

OpenStudy (whpalmer4):

Why not write it as \[\frac{(x^2-2x+1)^{x}}{x^2-2x+1}\]then you can apply L'Hopital a few times and Bob's your uncle...

OpenStudy (freckles):

only nee one round of l'hospital with my method I believe

OpenStudy (freckles):

\[\lim_{x \rightarrow 1^+}\ln[(x-1)^{2(x-1)]} \\ =\lim_{x \rightarrow 1^+} 2(x-1) \ln(x-1) \\ =2\lim_{x \rightarrow 1^+} \frac{\ln(x-1)}{\frac{1}{x-1}}\] apply l'hospital then go back and do e^(L) L being the answer to the above limit

OpenStudy (whpalmer4):

lots of folks learn about L'Hopital before they know the derivative of a log...my approach only requires chain rule and polynomial derivatives.

OpenStudy (freckles):

he found this limit earlier actually http://openstudy.com/users/freckles#/updates/56aad602e4b0c7c29515a067 I know you didn't know that though

OpenStudy (whpalmer4):

both approaches work, of course. all a matter of whether you prefer to pull the rabbit out of the hat at the beginning or the end of the trick :-)

OpenStudy (freckles):

I'm trying to figure out how to differentiate (x^2-2x+1)^x without knowing how to differentiate log

OpenStudy (amenah8):

thank you both! figured it out